OpenStudy (anonymous):

Find the length of the arc on a circle of radius r intercepted by a central angle θ. Round answer to two decimal places. r = 8 meters, θ = 50°

OpenStudy (anonymous):

Length of arc = radius of circle multiply by the angle in radians. Therefore, convert the angle first to radians through simple proportions. 180 degrees is equivalent to pi in radians Therefore, \[50 degrees \rightarrow \frac{ \pi }{ 180 } \times 50 = 0.873\] Therefore, \[Length of arc = 8 \times 0.873 = 6.98m\]

OpenStudy (anonymous):

THANK YOU sooo much :)

OpenStudy (anonymous):

my pleasure